# Cross-Sectional Dynamics Under Network Structure: Theory and Macroeconomic Applications

### v6

@@ 4.2 Application-Setup, Data & Estimation
I compute log differences of monthly industrial production (IP) indices across 23 manufacturing and mining sectors in the US economy provided by the Federal Reserve Board.The indices are available from January 2005 through August 2022. To eliminate seasonal patterns, I regress each series on month-dummies, take the residuals and add back the mean. To construct $A$, I use annual data on input-output (IO) matrices provided by the Bureau of Economic Analysis (BEA). I take the input-output matrix for 2010. Following the theory in Section 4.1, links $a_{ij}$ are calibrated as
@@
For the specification with lagged IOC, I consider $p=\{1,2,3,4,5,6\}$ and $q\in\{1,2,3\}$, implying network interaction frequencies of a month, two weeks and 10 days. Under contemporaneous IOC, I take $q=1$, as it already refers to the limit case of an infinitely high frequency of network interactions. In both cases, I consider Normal AR(1) processes for idiosyncratic and aggregate TFP: $(1-\rho_{a}L)e^{a}_{\tau}=\varepsilon^{a}_{\tau}\sim N(0,\sigma^{2}_{a})$ and $(1-\rho_{i}L)e_{i\tau}=\varepsilon_{i\tau}\sim N(0,\sigma^{2}_{i})$. To separately identify both TFP processes, I normalize $\sigma^{2}_{a}=1$ and $\delta_{1}=1$, re-defining $\delta$. To accommodate the restrictions $\alpha_{l}\geq 0 \;\forall\;l$ and $\sum_{l=1}^{p}\alpha_{l}=1$, I drop $\alpha_{p}$ from $\alpha$ and impose the domain restrictions $\alpha_{l}\in[0,1]$ for l=1:p1l=1:p-1 and $\sum_{l=1}^{p-1}\alpha_{l}\leq 1$. Under lagged IOC, this yields p1+4np-1+4n unknown parameters: $\theta=(\alpha^{\prime},\gamma^{\prime},\delta^{\prime},\rho_{a},\rho^{\prime},\sigma^{2\prime})^{\prime}$, where $\rho$ and $\sigma^{2}$ stack $\{\rho_{i}\}_{i=1}^{n}$ and $\{\sigma^{2}_{i}\}_{i=1}^{n}$, respectively. Under contemporaneous IOC, $\alpha$ is dropped from $\theta$, leading to $4n$ parameters.494949 With $n=116$ sectors and T=211pT=211-p periods available for estimation, this yields an observations-to-parameters ratio of 52.5 under $p=1$ and 50.7 under $p=6$.
@@ 4.3 Results
Although stronger effects tend to take more time to realize, there is no clear relationship between the strength and timing of the response of aggregate output to sectoral TFP shocks. This is illustrated by the right panel of Fig. 6. For example, a TFP shock in the food and beverage sector has a similarly small long-term effect on aggregate output as a shock to primary metals, yet the latter materializes much more sluggishly. One month after the TFP shock in the primary metals sector, a similar fraction of the long-term effect on aggregate output has materialized as in the case of a TFP shock to mining support activities, yet the latter are estimated to lead to a stronger long-term effect.
@@
In Section 4, a process $y_{t}$ is driven by an observed network, and we quantify how network effects materialize over time. In this section, I forecast a rather high-dimensional series $y_{t}$ and, for the most part, I assume that no network data is available. When $(\alpha,A)$ are jointly estimated and the estimation of $A$ is regularized, the NVAR becomes useful as a dimensionality-reduction technique. It rationalizes the dynamic comovement among all $\{y_{it}\}_{i=1:n}$ by the dynamic innovation transmission along a few bilateral links among units $i=1:n$.
In Section 5.1, I discuss the potential of the NVAR to reduce dimensionality, building on Section 3.2. Section 5.2 then sets up the application to forecast cross-contry macroeconomic aggregates, and Section 5.3 presents the results. Details are in Appendix D.
@@ 5.1 NVAR-Estimation for Dimensionality-Reduction
Even for intermediate $n$, an unrestricted VAR($p$) poorly forecasts $y_{t}\in\mathbb{R}^{n}$. Relative to it, under $p>1$, the NVAR($p,1$) reduces the number of parameters in the autoregressive matrices from $pn^{2}$ to $n^{2}+p-1$, owing to the assumption that innovations transmit cross-sectionally only via bilateral links; rather than freely relating $y_{it}$ to $\{y_{j,t-l}\}_{j=1:n,l=1:p}$, it uses the $n$ links $\{a_{ij}\}_{i=1:n}$ to compress this information to $x_{i,t-l}=\sum_{j=1}^{n}a_{ij}y_{j,t-l}$, and in turn the p1p-1 free parameters in $\alpha$ determine the importance of different lags of $x_{it}$ for the dynamics of $y_{it}$.515151 Even though $\mathbb{E}_{t-1}[y_{it}]$ contains the same linear combinations of $\{y_{j,t-l}\}_{j=1:n}$ at all lags $l=1:p$, dynamics at higher horizons $h$ are driven by different linear combinations, since higher-order network connections accumulate over time (see Proposition 1). In addition, the estimation of $A$ can be regularized, motivated by the sparseness of real world-networks across a wide range of applications and by the fact that the dynamic comovement of two series $y_{it}$ and $y_{jt}$ can be captured with higher-order connections between $i$ and $j$ without a direct link between them (see Propositions 1 and 2). In this case, the NVAR leverages both approaches available to address the large parameter problem in the Wold representation: it reduces the number of parameters and applies shrinkage priors (regularization) (see Geweke (1984)).
Section 3.2 discussed the estimation of $A$ under a Normal prior (L2-penalty) and under an Exponential prior (L1-penalty and restricting $a_{ij}\geq 0$). In the following, these two approaches are labeled NVAR-R and NVAR-L. The full posterior is obtained by iteratively drawing from the conditional posteriors of $\alpha|A$ and $A|\alpha$, while the posterior mode – the Ridge-/Lasso-regularized LS-/ML-estimator – is obtained by iterating on the conditional posterior modes.
When the NVAR-R and the NVAR-L are applied for predictive purposes, how to select the hyperparameter $\lambda_{a}$ that controls the degree of shrinkage becomes particularly important. Following Giannone et al. (2015), a standard approach is to set it to its marginal posterior mode (MPM) obtained under a Uniform prior, which maximizes the marginal data density (MDD) and, hence, the one-step ahead predictive ability. Under NVAR-R and NVAR-L, respectively, we obtain the conditional posteriors
@@
While this approach is grounded well in theory, it is computationally expensive, as it requires obtaining the full posterior.525252 This is particularly disadvantageous if only the posterior mode estimator is of interest. Even if the full posterior of $(\alpha,A)$ evaluated at the optimal $\lambda_{a}$ is desired, however, it is a drawback, as it requires conducting posterior sampling twice. The disadvantage disappears only if one is interested in the marginal posterior of $(\alpha,A)$ under a Uniform prior for $\lambda_{a}$ (model-averaging rather than -selection). For NVAR-L, a heuristic approach is to maximize the Bayesian Information Criterion (BIC) suggested by Zou et al. (2007) for other Lasso-applications. This involves counting the number of non-zero elements in $\hat{A}(\lambda_{a})$. For NVAR-R, an analogous criterion is the conditional MDD $p(Y|\lambda_{a},B,\alpha,\Sigma)$, which is derived in Section D.1 and can be maximized when evaluated at $\hat{\alpha}$ (and $\hat{\Sigma}$).
@@ 5.2 Forecasting-Setup
where $u_{t}\sim N(0,I_{n})$, $\eta_{t}\sim N(0,I_{r})$ and $\Lambda_{1:r,\cdot}$ is lower-triangular with positive diagonal elements. The equivalence result in Proposition 4 suggests that, relative to the DFM, the NVAR restricts factor dynamics, but allows the linear combinations of predictors to vary more flexibly across units and acommodates sparse factors as (linear combinations of) locally important units in the network. Therefore, the NVAR is expected to capture cross-sectional dynamics in finite samples better than the factor model in the presence of many sparse factors or, equivalently, in case of a sparse, yet close-to-full-rank network adjacency matrix. Intuitively, this is the case of dynamics driven by many bilateral links rather than a few influential units. More generally, Proposition 4 suggests the NVAR could improve upon the poor forecasting performance of factor models under a high dispersion of factor loadings across series (see Boivin and Ng (2006)).535353 This notably includes the case of sparse factors, i.e. loading-vectors with zero and non-zero entries. As Boivin and Ng (2006) point out, selecting the number of factors separately for each series is a poor remedy: series that depend on less dominant factors are still poorly forecasted, as including more estimated factors adds noise to the forecasts.
To investigate these hypotheses, I use the NVAR and DFM to forecast monthly industrial production (IP) growth, monthly CPI inflation and quarterly real GDP growth across OECD-member countries, -applicants and -partners.545454 As of August 2024, this includes 45 countries. On top of 34 OECD members, there are 8 applicants (Argentina, Brazil, Bulgaria, Croatia, Indonesia, Peru, Romania and Thailand) and 3 partners (China, India, South Africa). The series are obtained from the IMF’s IFS database, and YoY growth rates are used. I limit attention to the pre-COVID periods 2001:M1 - 2019:M12 and 2001:Q1 - 2019:Q4, respectively, and I delete countries with missings or more than two consecutive constants. This yields datasets with $n=42$, $n=34$ and $n=39$ observations. The series are de-seasonalized by subtracting fitted values from a regression on period-dummies, and the resulting series are standardized.
@@
I limit the analysis to point forecasts obtained under the posterior mode of $(\alpha,A)$ in the NVAR-R and NVAR-L as presented in Section 3.2 – the Ridge-/Lasso-regularized ML-/LS-estimator –, and the posterior mode of $(\Lambda,\Phi)$, $\Phi=[\Phi_{1},...,\Phi_{p}]^{\prime}$ under a Uniform prior in the DFM – the ML-/LS-estimator (see Section D.3). For the DFM, I consider $p=1:4$ lags and $r=10$ factors. In turn, I select the ex-post best-performing specification that minimizes the cumulative MSE over the first three horizons, and I compare all NVAR specifications against this benchmark. This renders the assessment of the NVAR’s suitability for forecasting independent of methods to select the number of factors ex-ante. For the NVAR, I also consider $p=1:4$ lags, and I mostly focus on $q=1$ – the case referred to by Proposition 4 . In both NVAR models, I set $\Sigma=I$, and, unless otherwise stated, I shrink $A$ to $B=0$. The degree of shrinkage, as embodied by $\lambda_{a}$, is chosen based on two methods. The first takes the MPM and, therefore, maximizes the MDD exactly, while the second approximates this choice by using grid-search to maximize the BIC (in case of the NVAR-L) or the conditional MDD (for the NVAR-R).
@@ 5.3 Results
The results in Fig. 7 are obtained using the respective posterior mode, i.e. frequentist point estimator. As shown in Fig. A-1, the factor models yield similar performances even for forecasts at the posterior mean or the posterior mean forecasts. Selecting the best DFM using these alternative forecasting-types does not change any of the above numbers by more than two percentage points. For NVAR-L, the normalization $\alpha_{1}=1$ is applied, whereas for NVAR-R, $\alpha||_{1}=1$ is imposed.565656 Under the former normalization, NVAR-R yields poor performance, as the free parameters in $\alpha$ are increased to extremely high values, while all elements in $A$ are shrunk accordingly. This is likely because, for highly correlated series with $\mathbb{E}[y_{t}y_{t-1}^{\prime}]\approx\mathbb{E}[y_{t}y_{t-2}^{\prime}]$, $\alpha_{-1}$ is weakly identified, as e.g. $Ay_{t-1}+\alpha_{2}Ay_{t-2}\approx(1+\alpha_{2})Ay_{t-1}$. This issue does not occur for NVAR-L, likely because it shrinks numerous elements of $A$ all the way to zero. The issue does not occur either when estimating the full posterior of NVAR-R or NVAR-L. The computational time needed to obtain the posterior modes of the DFM and the NVAR for a given $\lambda_{a}$ are of a similar order of magnitude and amount to 5-10 seconds. However, this burden is considerably increased by the need to re-compute the posterior mode for many different $\lambda_{a}$ in the search for the value that maximizes MDD or BIC and the need to compute the full posterior distribution in the search for the MPM of $\lambda_{a}$. [^57] For NVAR-R, this yields about 50-100s in the former case and about 10min in the latter case. These times are doubled for NVAR-L.

### v7

@@ 4.2 Application-Setup, Data & Estimation
1818todo: 18need to discuss aggregation weights1919todo: 19need to do a more careful decomposition analysis; see latex comment
@@
I compute log differences of monthly industrial production (IP) indices across 23 manufacturing and mining sectors in the US economy provided by the Federal Reserve Board.The indices are available from January 2005 through August 2022. 2020todo: 20sure? not later? To eliminate seasonal patterns, I regress each series on month-dummies, take the residuals and add back the mean. To construct $A$, I use annual data on input-output (IO) matrices provided by the Bureau of Economic Analysis (BEA). I take the input-output matrix for 2010. Following the theory in Section 4.1, links $a_{ij}$ are calibrated as
@@
For the specification with lagged IOC, I consider $p=\{1,2,3,4,5,6\}$ and $q\in\{1,2,3\}$, implying network interaction frequencies of a month, two weeks and 10 days. Under contemporaneous IOC, I take $q=1$, as it already refers to the limit case of an infinitely high frequency of network interactions. In both cases, I consider Normal AR(1) processes for idiosyncratic and aggregate TFP: $(1-\rho_{a}L)e^{a}_{\tau}=\varepsilon^{a}_{\tau}\sim N(0,\sigma^{2}_{a})$ and $(1-\rho_{i}L)e_{i\tau}=\varepsilon_{i\tau}\sim N(0,\sigma^{2}_{i})$. To separately identify both TFP processes, I normalize $\sigma^{2}_{a}=1$ and $\delta_{1}=1$, re-defining $\delta$. To accommodate the restrictions $\alpha_{l}\geq 0 \;\forall\;l$ and $\sum_{l=1}^{p}\alpha_{l}=1$, I drop $\alpha_{p}$ from $\alpha$ and impose the domain restrictions $\alpha_{l}\in[0,1]$ for l=1:p1l=1:p-1 and $\sum_{l=1}^{p-1}\alpha_{l}\leq 1$. Under lagged IOC, this yields p1+4np-1+4n unknown parameters: $\theta=(\alpha^{\prime},\gamma^{\prime},\delta^{\prime},\rho_{a},\rho^{\prime},\sigma^{2\prime})^{\prime}$, where $\rho$ and $\sigma^{2}$ stack $\{\rho_{i}\}_{i=1}^{n}$ and $\{\sigma^{2}_{i}\}_{i=1}^{n}$, respectively. Under contemporaneous IOC, $\alpha$ is dropped from $\theta$, leading to $4n$ parameters. [^39] With $n=116$ sectors and T=211pT=211-p periods available for estimation, this yields an observations-to-parameters ratio of 52.5 under $p=1$ and 50.7 under $p=6$.
@@ 4.3 Results
2323todo: 23put IRFs on RHS top of each other2424todo: 24put in footnote: this is like mediation analysis in Dufour & Wang (2023); they do it by variable, I do it by connection order.
@@
Although stronger effects tend to take more time to realize, there is no clear relationship between the strength and timing of the response of aggregate output to sectoral TFP shocks. This is illustrated by the right panel of Fig. 6. For example, a TFP shock in the food and beverage sector has a similar long-term effect on aggregate output as a shock to primary metals, yet the latter materializes much more sluggishly. One month after the TFP shock in the primary metals sector, a similar fraction of the long-term effect on aggregate output has materialized as in the case of a TFP shock to mining support activities, yet the latter are estimated to lead to a stronger long-term effect.
@@
In Section 4, a process $y_{t}$ is driven by an observed network, and we quantify how network effects materialize over time. In this section, I forecast a rather high-dimensional series $y_{t}$ and, for the most part, I assume that no network data is available.2525todo: 25”for the most part”… When $(\alpha,A)$ are jointly estimated and the estimation of $A$ is regularized, the NVAR becomes useful as a dimensionality-reduction technique. It rationalizes the dynamic comovement among all $\{y_{it}\}_{i=1:n}$ by the dynamic innovation transmission along a few bilateral links among units $i=1:n$.
In Section 5.1, I discuss the potential of the NVAR to reduce dimensionality, building on Section 3.2. Section 5.2 then sets up the application to forecast macroeconomic aggregates across countries, and Section 5.3 presents the results. Details are in Appendix D.
@@ 5.1 NVAR-Estimation for Dimensionality-Reduction
Even for intermediate $n$, an unrestricted VAR($p$) poorly forecasts $y_{t}\in\mathbb{R}^{n}$.2626todo: 26cite.. Relative to it, under $p>1$, the NVAR($p,1$) reduces the number of parameters in the autoregressive matrices from $pn^{2}$ to $n^{2}+p-1$, owing to the assumption that innovations transmit cross-sectionally only via bilateral links; rather than freely relating $y_{it}$ to $\{y_{j,t-l}\}_{j=1:n,l=1:p}$, it uses the $n$ links $\{a_{ij}\}_{i=1:n}$ to compress this information to $x_{i,t-l}=\sum_{j=1}^{n}a_{ij}y_{j,t-l}$, and in turn the p1p-1 free parameters in $\alpha$ determine the importance of different lags of $x_{it}$ for the dynamics of $y_{it}$. [^41] Even though $\mathbb{E}_{t-1}[y_{it}]$ contains the same linear combinations of $\{y_{j,t-l}\}_{j=1:n}$ at all lags $l=1:p$, dynamics at higher horizons $h$ are driven by different linear combinations, since higher-order network connections accumulate over time (see Proposition 1). If, in addition, $A$ is estimated parsimoniously, then the NVAR leverages both approaches available to address the large parameter problem in the Wold representation (Geweke, 1984): it reduces the number of parameters and applies shrinkage priors (regularization). Regularization of $A$ is motivated by the sparseness of real world-networks across a wide range of applications2727todo: 27cite and by the fact that the dynamic comovement of two series $y_{it}$ and $y_{jt}$ can be captured with higher-order connections between $i$ and $j$ without a direct link between them (see Propositions 1 and 2).
Section 3.2 discussed the estimation of $A$ under a Normal prior (L2-penalty) and under an Exponential prior (L1-penalty and restricting $a_{ij}\geq 0$). In the following, these two approaches are labeled NVAR-R and NVAR-L. The joint posterior of $(\alpha,A)$ is obtained by iteratively drawing from the conditional posteriors of $\alpha|A$ and $A|\alpha$, the joint posterior mode – the Ridge-/Lasso-regularized LS-/ML-estimator – by iterating on the conditional posterior modes.
When the NVAR-R and NVAR-L are applied for predictive purposes, selecting the hyperparameter $\lambda_{a}$ – i.e. the degree of shrinkage – becomes particularly important. Following Giannone et al. (2015), setting a hyperparameter to its marginal posterior mode (MPM) under a Uniform hyperprior maximizes the marginal data density (MDD) and, hence, the one-step ahead predictive ability.2828todo: 28optimizes … under a quadratic loss function? Under NVAR-R and NVAR-L, respectively, we obtain the conditional posteriors
@@
While well-grounded in theory, the approach of Giannone et al. (2015) is computationally expensive in the present environment, as it requires obtaining the full posterior. [^42] This is particularly disadvantageous if only the posterior mode estimator is of interest. Even if the full posterior of $(\alpha,A)$ evaluated at the optimal $\lambda_{a}$ is desired, however, it is a drawback, as it requires conducting posterior sampling twice. The disadvantage disappears only if one is interested in the marginal posterior of $(\alpha,A)$ under a Uniform prior for $\lambda_{a}$ (model-averaging rather than -selection). For NVAR-L, a heuristic approach is to maximize the Bayesian Information Criterion (BIC) suggested by Zou et al. (2007) for other Lasso-applications. This involves counting the number of non-zero elements in $\hat{A}(\lambda_{a})$. For NVAR-R, an analogous criterion is the conditional MDD $p(Y|\lambda_{a},B,\alpha,\Sigma)$, which is derived in Section D.1 and can be maximized when evaluated at $\hat{\alpha}$ (and $\hat{\Sigma}$).
2929todo: 29see commented-out text on interpretability and other
@@ 5.2 Forecasting-Setup
where $u_{t}\sim N(0,I_{n})$, $\eta_{t}\sim N(0,I_{r})$ and $\Lambda_{1:r,\cdot}$ is lower-triangular with positive diagonal elements. The equivalence result in Proposition 4 suggests that, relative to the DFM, the NVAR restricts factor dynamics, but allows the linear combinations of predictors to vary more flexibly across units and acommodates sparse factors as (linear combinations of) locally important units in the network. Therefore, the NVAR is expected to capture cross-sectional dynamics in finite samples better than the factor model in the presence of many sparse factors or, equivalently, in case of a sparse, yet close-to-full-rank network adjacency matrix. Intuitively, this is the case of dynamics driven by many bilateral links rather than a few influential units. More generally, Proposition 4 suggests the NVAR could improve upon the poor forecasting performance of factor models under a high dispersion of factor loadings across series (see Boivin and Ng (2006)). [^43] This notably includes the case of sparse factors, i.e. loading-vectors with zero and non-zero entries. As Boivin and Ng (2006) point out, selecting the number of factors separately for each series is a poor remedy: series that depend on less dominant factors are still poorly forecasted, as including more estimated factors adds noise to the forecasts.
3030todo: 30Can we show that in results? note that it holds not unit-by unit, as I thoguht initially (see commented-out text here)
To investigate these hypotheses, I use the NVAR and DFM to forecast monthly industrial production (IP) growth, monthly CPI inflation and quarterly real GDP growth across OECD-member countries, -applicants and -partners. [^44] As of August 2024, this includes 45 countries. On top of 34 OECD members, there are 8 applicants (Argentina, Brazil, Bulgaria, Croatia, Indonesia, Peru, Romania and Thailand) and 3 partners (China, India, South Africa). The series are obtained from the IMF’s IFS database, and YoY growth rates are used. I limit attention to the pre-COVID periods 2001:M1 - 2019:M12 and 2001:Q1 - 2019:Q4, respectively, and I delete countries with missings or more than two consecutive constants. This yields datasets with $n=42$, $n=34$ and $n=39$ observations. The series are de-seasonalized by subtracting fitted values from a regression on period-dummies, and the resulting series are standardized.
@@
I limit the analysis to point forecasts obtained under the posterior mode of $(\alpha,A)$ in the NVAR-R and NVAR-L as presented in Section 3.2 – the Ridge-/Lasso-regularized ML-/LS-estimator –, and the posterior mode of $(\Lambda,\Phi)$, $\Phi=[\Phi_{1},...,\Phi_{p}]^{\prime}$ under a Uniform prior in the DFM – the ML-/LS-estimator (see Section D.3).3131todo: 31and orthogonal factors! For the DFM, I consider $p=1:4$ lags and $r=10$ factors. In turn, I select the ex-post best-performing specification that minimizes the cumulative MSE over the first three horizons, and I compare all NVAR specifications against this benchmark. This renders the assessment of the NVAR’s suitability for forecasting independent of methods to select the number of factors ex-ante. For the NVAR, I also consider $p=1:4$ lags, and I mostly focus on $q=1$ – the case referred to by Proposition 4 3232todo: 32mostly…. In both NVAR models, I set $\Sigma=I$, and, unless otherwise stated, I shrink $A$ to $B=0$. The degree of shrinkage, as embodied by $\lambda_{a}$, is chosen based on two methods. The first takes the MPM and, therefore, maximizes the MDD exactly, while the second approximates this choice by using grid-search to maximize the BIC (in case of the NVAR-L) or the conditional MDD (for the NVAR-R).
@@ 5.3 Results
The results in Fig. 7 are obtained using the respective posterior mode, i.e. frequentist point estimator. As shown in Fig. A-1, the factor models yield similar performances even for forecasts at the posterior mean or the posterior mean forecasts. Selecting the best DFM using these alternative forecasting-types does not change any of the above numbers by more than two percentage points. For NVAR-L, the normalization $\alpha_{1}=1$ is applied, whereas for NVAR-R, $\alpha||_{1}=1$ is imposed. [^46] Under the former normalization, NVAR-R yields poor performance, as the free parameters in $\alpha$ are increased to extremely high values, while all elements in $A$ are shrunk accordingly. This is likely because, for highly correlated series with $\mathbb{E}[y_{t}y_{t-1}^{\prime}]\approx\mathbb{E}[y_{t}y_{t-2}^{\prime}]$, $\alpha_{-1}$ is weakly identified, as e.g. $Ay_{t-1}+\alpha_{2}Ay_{t-2}\approx(1+\alpha_{2})Ay_{t-1}$. This issue does not occur for NVAR-L, likely because it shrinks numerous elements of $A$ all the way to zero. The issue does not occur either when estimating the full posterior of NVAR-R or NVAR-L. The computational time needed to obtain the posterior modes of the DFM and the NVAR for a given $\lambda_{a}$ are of a similar order of magnitude and amount to 5-10 seconds. However, this burden is considerably increased by the need to re-compute the posterior mode for many different $\lambda_{a}$ in the search for the value that maximizes MDD or BIC and the need to compute the full posterior distribution in the search for the MPM of $\lambda_{a}$. [^47] For NVAR-R, this yields about 50-100s in the former case and about 10min in the latter case. These times are doubled for NVAR-L.
@@
3333todo: 33see commented-out section on $B$-selection here and whole appendix-section3434todo: 34see commented-out section on $q\neq 1$ here